# Agent Skill Acquisition for Large Language Models via CycleQD

### v1

@@ 1 Introduction
In response to these challenges, we propose CycleQD, a novel framework that leverages the Quality Diversity (QD) method (Pugh et al., 2016), specifically through a cyclic adaptation of the MAP-Elites algorithm (Mouret & Clune, 2015). CycleQD features a dynamic cycle where each agentic skill’s performance metric (e.g., pass@1 for coding tasks, success rate for Operation System (OS) and Database (DB) tasks) is optimized in isolation, with other skills represented as behavioral characteristics (BCs). Unlike conventional crossover operations that randomly swap the parameters between two genes, CycleQD employs a model merging based crossover operation (Akiba et al., 2024), which is essential for transferring skills from specialized experts to a new, cohesive model. Furthermore, we also introduce a singular value decomposition (SVD)-based mutation method to extrapolate model capabilities while preventing overfitting. Because CycleQD uses task performance as its descriptors, it can be viewed as a simplified version of Multi-Objective QD (MOQD) (Pierrot et al., 2022) that optimizes in a coordinate ascent fashion, without the need for complex hyper-volume calculations when the BC dimension increases. See Figure 1 for an overview of CycleQD.
@@
2 Related works
@@ 3 Methods
  1. 1.
    Alternating Quality and BCs: Pugh et al. (2016) emphasizes the importance of well-aligned BCs with the quality (i.e., task metric that needs to be optimized). Traditional systems typically set the design at the onset and do not alter them throughout the QD process. In contrast, CycleQD uses task metrics as both quality and BCs, ensuring alignment and dynamically alternating them during the QD process (lines 6-7 in Algorithm 1).
  2. 2.
    Model Merging as Crossover: Unlike existing systems that optimize model parameters directly, CycleQD leverages a model merging algorithm as a crossover operation, replacing the heuristic, hand-designed rules often seen in practice (line 12 in Algorithm 1). In CycleQD, we fine-tune expert agents, each of which only needs to specialize in one task, and use them as the seed models to initialize the model archives (lines 2-4 in Algorithm 1).
  3. 3.
    SVD-based Mutation: Traditional mutation operations in genetic algorithms like MAP-Elites typically introduce random perturbations from a pre-defined distribution to explore new regions. In contrast, CycleQD utilizes perturbations aligned with the principal components of the models’ parameters matrices, facilitating exploration outside the convex regions formed by the parent models while avoiding overfitting (line 13 in Algorithm 1).
@@ 3.1 Alternating quality and behavior characteristics
We employ task performance metrics as our BCs, imbuing them with concrete performance implications. This allows for the design of an Elite sampling algorithm aimed to expand the Pareto frontier. This Elite sampling algorithm is specifically designed for the CycleQD setup, where both quality and BCs reflect task performance. In this context, it is more effective to prioritize models that push the frontier of higher performance. This frontier is defined by the models achieving high performance across these BCs. Concretely, when the $i$-th archive is active, a model $j$ inside it is sampled with probability $P_{j}=\frac{\gamma_{j}}{\sum_{n=1}^{N}{\gamma_{n}}}$ where $N$ is the number of models in this archive and $\gamma_{j}$ is calculated as follows:
@@ 3.2 Model merging based crossover
Training an LLM to specialize in a single task does not suffer from problems such as the data ratio tuning, the complex combinations of objective functions, and etc. It is therefore straightforward to initially train a set of single-task experts using conventional fine-tuning methods, and then use model merging algorithms to develop more capable agents. In CycleQD, our model merging crossover operator employs the parameter space merging scheme described in Akiba et al. (2024), which creates a new model by merging task vectors at the model level (Ilharco et al., 2022). Specifically, for a pre-trained base LLM with parameters $\theta_{\mathrm{base}}\in\mathbb{R}^{d}$ and its fine-tuned LLM with parameters $\theta\in\mathbb{R}^{d}$, we define a task vector as $\tau=\theta-\theta_{\mathrm{base}}$. The crossover operator then generates a model’s parameters by combining these task vectors: $\theta_{\mathrm{child}}=\theta_{\mathrm{base}}+\big{(}\omega_{1}/(\omega_{1}+\omega_{2})\big{)}\tau_{p_{1}}+\big{(}\omega_{2}/(\omega_{1}+\omega_{2})\big{)}\tau_{p_{2}}$, where $\tau_{p_{1}}$ and $\tau_{p_{2}}$ are the parents’ task vectors. Here, $\omega_{1}$ and $\omega_{2}$ are i.i.d samples from $\mathcal{N}(\mu,\sigma^{2})$, and $(\mu,\sigma)$ are predetermined hyper-parameters that remain fixed during the experiments. We normalize $(\tau_{p_{1}},\tau_{p_{2}})$’s mixing coefficients to ensure the merged weights do not become outliers and cause problems for the downstream layers.
@@ 3.3 SVD-based mutation
Intuitively, if task vectors are representative of the agent skills that each fine-tuned LLM has mastered, their components are the building blocks and are therefore proxies in a finer granularity. As a result, our SVD-based mutation allows finer control of $\theta_{\mathrm{child}}$ in the fundamental skill building blocks and reduces the risk of overfitting. See Figure 2 for an illustration.
@@ 3.4 Model aggregation
CycleQD maintains $K$ archives during the optimization process, LLMs in each of these archives are trained to acquire one of the $K$ agent skills. In the end, CycleQD returns hundreds or even thousands of LLM agents with various specialties and characteristics. These LLM-based agents are versatile and can facilitate multi-agent research in a diverse set of directions (we discuss this topic in Section 5). On the other hand, an aggregated LLM agent that possesses the top skills acquired by the best models in each archive is useful on its own and is important for the evaluation of CycleQD.
Our model aggregation algorithm is straightforward, and can be summarized as: $\theta_{\mathrm{agg}}=\theta_{\mathrm{base}}+\sum_{k=1}^{K}{\beta_{k}\tau_{k}}$, where $\tau_{k}$ is the task vector from the elite model in the $k$-th archive, with ties resolved by selecting the most recent model. $\beta_{k}=\exp(f_{k})/\sum_{i=1}^{K}{\exp(f_{i})}$ (i.e., softmax function) is the mixing coefficient calculated from the elite model’s task performance $f_{k}$ in the $k$-th archive. In this paper, all experimental results from CycleQD are produced with this aggregated model $\theta_{\mathrm{agg}}$.
@@ 4 Experiments
We evaluate CycleQD extensively and conduct analysis to answer the following questions:
  • Does CycleQD allow LLMs to effectively acquire various agent skills? If so, how good is it when compared with baseline methods?
  • What are the critical design choices that lead to its effectiveness?
  • What are the benefits and limitations of CycleQD?
@@ 4.1.3 Baselines
Merging Based Models: These are models 7-9 in Table 1. We first introduce a naive merging method where the merged model is produced by taking an average of the three experts’ task vectors (model 7). Similar to this baseline, we introduce two additional learning based model merging methods where, instead of taking an average, the coefficients of a linear combination of these task vectors is learned through gradient descent on policy gradients (model 8), and CMA-ES, an evolutionary algorithm on the raw rewards (model 9) separately.
@@ 4.1.4 Results
Comparison with Merging Based Methods: CycleQD also outperforms conventional model merging based methods (models 7-9), and we conjecture that this is due to two reasons. First, these methods lack a mutation operation, which we believe is crucial for enabling the merged model to escape the “convex region” created by the seed models. For example, model 9, while having the best average performance in its category, scores lower on all three tasks compared to the experts. In contrast, CycleQD significantly surpasses expert performance in coding and OS tasks. Second, we suspect that the absence of a mechanism to rehabilitate temporarily underperforming models contributes to this discrepancy. Given the lengthy computation pipelines in LLMs, which include numerous layers, even a single malfunctioning layer can degrade overall performance. Unlike QD, the baselines do not have a way to recover these layers, potentially promising models are prematurely discarded.
@@ 4.1.5 Ablation studies
Sampling Methods: Finally, we demonstrate the importance of Elite sampling in CycleQD in trial 4, which has the identical settings as model 10 in Table 1, and gives the best score in the ablation studies. Elite sampling is effective because it prioritizes merging high-performing models, the result of which helps expand the Pareto frontier in each archive more effectively.
@@ 4.1.6 Skill generalization and language capabilities
In addition to agent skills, we evaluate our model on out-of-distribution coding tasks and language understanding across six categories, please see Appendix A.1.5 for detailed task descriptions. Table 3 shows the results across these tasks, where we normalize the scores against the Llama3-8B-Instruct base model. On average, CycleQD outperforms the three fine-tuned experts, providing better overall generalization across tasks. Our method generalizes well on the coding tasks while retraining performance on language tasks. In contrast, the Coding expert, while delivering competitive scores on the coding tasks, shows significantly lower performance on specific tasks in the Reasoning category. This sharp decline in performance suggests the occurrence of catastrophic forgetting during fine-tuning, where the model loses its ability to generalize across tasks outside its specialization. These findings once again underscore the efficacy of CycleQD in achieving superior and more consistent performance across diverse tasks compared to traditional fine-tuning approaches.
@@ 4.2.1 Task description
In addition to its applications for LLMs, CycleQD serves as a versatile method for integrating expert models across various data modalities beyond text. In this experiment, we extend CycleQD to the fusion of multiple Segment Anything Models (SAM), which are state-of-the-art computer vision models designed for image segmentation tasks. Specifically, our objective is to merge pairs of SAM models, A and B, to create models whose capabilities encompass the skill sets of both A and B.
@@ Acknowledgements
The authors would like to thank Robert Lange for providing valuable discussions and feedback while drafting the text, and Kou Misaki for providing infrastructure support. This paper is based on results obtained from a project, JPNP20017, subsidized by the New Energy and Industrial Technology Development Organization (NEDO).
@@ A.1.3 CycleQD hyper-parameters
We set $\alpha_{\mathrm{low}}=0.5$ and $\alpha_{\mathrm{high}}=0.8$ in Elite sampling, $\mu=1.0$ and $\sigma=0.03$ in model merging-based crossover, and $w_{\text{max}}=0.3$ in our SVD-based mutations.

### v2

@@ 1 Introduction
In response to these challenges, we propose CycleQD, a novel framework that leverages the Quality Diversity (QD) method (Pugh et al., 2016), specifically through a cyclic adaptation of the MAP-Elites algorithm (Mouret & Clune, 2015). CycleQD features a dynamic cycle where each agentic skill’s performance metric (e.g., pass@1 for coding tasks, success rate for Operation System (OS) and Database (DB) tasks) is optimized in isolation, with other skills represented as behavioral characteristics (BCs). Unlike conventional crossover operations that randomly swap the parameters between two genes, CycleQD employs a model merging based crossover operation (Akiba et al., 2024), which is essential for transferring skills from specialized experts to a new, cohesive model. Furthermore, we also introduce a singular value decomposition (SVD)-based mutation method to extrapolate model capabilities while preventing overfitting. Unlike gradient-based fine-tuning, CycleQD reduces the need for intricate design decisions and hyperparameter tuning, while inherently capturing diverse skills and behaviors. See Figure 1 for an overview of CycleQD.
@@
2 Preliminaries
Figure 2: MAP-Elites flow.
Evolutionary Strategies (ES) are optimization algorithms inspired by natural selection. They improve a population of candidate solutions through mutation, crossover (i.e., recombination), and selection, guided by a fitness function. By maintaining diversity, ES avoids premature convergence and is well-suited for QD optimization, which aims to generate diverse high-performing solutions rather than a single global optimum.
MAP-Elites is a key algorithm within the QD paradigm, designed to optimize for both performance and diversity by maintaining an archive of elite solutions across a discretized space of behavioral characteristics (BCs). Each cell in this archive corresponds to a unique combination of BCs, storing the best solution found for that cell. Figure 2 illustrates an optimization step: two parents are sampled from the archive, and a child solution is generated through crossover and mutation. The child’s fitness and BCs are then evaluated to determine its placement in the archive, replacing the current entry if it performs better.
@@ 3 Methods
Alternating Quality and BCs: Pugh et al. (2016) emphasizes the importance of well-aligned BCs with the quality (i.e., task metric that needs to be optimized). Traditional systems typically set the design at the onset and do not alter them throughout the QD process. In contrast, CycleQD uses task metrics as both quality and BCs, ensuring alignment and dynamically alternating them during the QD process (lines 6-7 in Algorithm 1). This step can be viewed as if the archive is rotated by 90 degrees after each generation before proceeding to the optimization, as is illustrated in Figure 1.
Model Merging as Crossover: Unlike existing systems that optimize model parameters directly, CycleQD leverages a model merging algorithm as a crossover operation, replacing the heuristic, hand-designed rules often seen in practice (line 12 in Algorithm 1). I.e., in Figure 2, our crossover operation merges the two parent models. In CycleQD, we fine-tune expert agents, each of which only needs to specialize in one task, and use them as the seed models to initialize the model archives (lines 2-4 in Algorithm 1).
SVD-based Mutation: Traditional mutation operations in genetic algorithms like MAP-Elites typically introduce random perturbations from a pre-defined distribution to explore new regions (i.e., the mutation operation in Figure 2). In contrast, CycleQD utilizes perturbations aligned with the principal components of the models’ parameters matrices, facilitating exploration outside the convex regions formed by the parent models while avoiding overfitting (line 13 in Algorithm 1).
CycleQD repeats these three steps in each generation, and the archives, originally only occupied by the expert models, accommodate more and more capable and diverse models, as is illustrated by Figure 6 in the Appendix.
@@ 3.1 Alternating quality and behavior characteristics
We employ task performance metrics as our BCs, imbuing them with concrete performance implications. This allows for the design of an Elite sampling algorithm aimed to expand the Pareto frontier. Our Elite sampling algorithm is inspired by Wang et al. (2023) which expand the Pareto frontier not only in quality but also within BCs. However, there are two major differences in CycleQD’s setup: (1) Both Q and BCs are task metrics and are treated equally in a cyclic process, (2) We aggregate Q and BC values into a single metric to drive the frontier towards higher performance. Specifically, this frontier is defined by the models achieving high performance across these BCs. Concretely, when the $i$-th archive is active, a model $j$ inside it is sampled with probability $P_{j}=\frac{\gamma_{j}}{\sum_{n=1}^{N}{\gamma_{n}}}$ where $N$ is the number of models in this archive and $\gamma_{j}$ is calculated as follows:
@@ 3.2 Model merging based crossover
Training an LLM to specialize in a single task does not suffer from problems such as the data ratio tuning, the complex combinations of objective functions, and etc. It is therefore straightforward to initially train a set of single-task experts using conventional fine-tuning methods, and then use model merging algorithms to develop more capable agents. In CycleQD, our model merging crossover operator employs the parameter space merging scheme described in Akiba et al. (2024), which creates a new model by merging task vectors at the model level (Ilharco et al., 2022). Specifically, for a pre-trained base LLM with parameters $\theta_{\mathrm{base}}\in\mathbb{R}^{d}$ and its fine-tuned LLM with parameters $\theta\in\mathbb{R}^{d}$, we define a task vector as $\tau=\theta-\theta_{\mathrm{base}}$. The crossover operator then generates a model’s parameters by combining these task vectors: $\theta_{\mathrm{child}}=\theta_{\mathrm{base}}+\big{(}\omega_{1}/(\omega_{1}+\omega_{2})\big{)}\tau_{p_{1}}+\big{(}\omega_{2}/(\omega_{1}+\omega_{2})\big{)}\tau_{p_{2}}$, where $\tau_{p_{1}}$ and $\tau_{p_{2}}$ are the parents’ task vectors. Here, $\omega_{1}$ and $\omega_{2}$ are i.i.d samples from $\mathcal{N}(\mu,\sigma^{2})$, and $(\mu,\sigma)$ are predetermined hyper-parameters that remain fixed during the experiments. These $\omega$’s do not need to be positive numbers. By allowing negative component weights, the merged model has more freedom in optimizing its task vectors, the process of which is automatically carried out via evolutionary search. We normalize $(\tau_{p_{1}},\tau_{p_{2}})$’s mixing coefficients to ensure the merged weights do not become outliers and cause problems for the downstream layers.
@@ 3.3 SVD-based mutation
This approach provides two key advantages. First, SVD focuses mutations on the fundamental building blocks of the task vector, which represent the essential components of agent skills. By constraining perturbations to these meaningful directions, we avoid the excessive exploration inherent in random perturbations, reducing the risk of overfitting. Second, the ability to manipulate individual components (i.e., $\omega$) offers finer control over $\theta_{\texttt{child}}$, enabling targeted modifications at a granular level. Figure 3 illustrates this process.
@@ 3.4 Model aggregation
CycleQD maintains $K$ archives during the optimization process, LLMs in each of these archives are trained to acquire one of the $K$ agent skills. In the end, CycleQD returns hundreds or even thousands of LLM agents with various specialties and characteristics. These LLM-based agents are versatile and can facilitate multi-agent research in a diverse set of directions (we discuss this topic in Section 6).
On the other hand, our goal in this work is to create a single model that achieves high performance across multiple tasks, and a model aggregation method is therefore necessary. Our model aggregation algorithm is straightforward, and can be summarized as: $\theta_{\mathrm{agg}}=\theta_{\mathrm{base}}+\sum_{k=1}^{K}{\beta_{k}\tau_{k}}$, where $\tau_{k}$ is the task vector from the elite model in the $k$-th archive, with ties resolved by selecting the most recent model. $\beta_{k}=\exp(f_{k})/\sum_{i=1}^{K}{\exp(f_{i})}$ (i.e., softmax function) is the mixing coefficient calculated from the elite model’s task performance $f_{k}$ in the $k$-th archive. In this paper, all experimental results from CycleQD are produced with this aggregated model $\theta_{\mathrm{agg}}$.
@@ 4.1.3 Baselines
Merging Based Models: These are models 7-10 in Table 1. We first introduce a naive merging method where the merged model is produced by taking an average of the three experts’ task vectors (model 7). Similar to this baseline, we introduce three additional learning based model merging methods where, instead of taking an average, the coefficients of a linear combination of these task vectors is learned through gradient descent on policy gradients (model 8), CMA-ES, an evolutionary algorithm on the raw rewards (model 9), and NSGA-II (Deb et al., 2002), another evolutionary algorithm that optimizes multiple objectives simultaneously (model 10).
@@ 4.1.4 Results
Comparison with Merging Based Methods: CycleQD also outperforms conventional model merging based methods (models 7-10), and we conjecture that this is due to three reasons. First, these methods lack a mutation operation, which we believe is crucial for enabling the merged model to escape the “convex region” created by the seed models. For example, model 9 scores lower on all three tasks compared to the experts. In contrast, CycleQD significantly surpasses expert performance in coding and OS tasks. Second, CycleQD’s cyclic optimization mechanism incorporates elements resembling non-dominant sorting in NSGA-II (model 10). However, unlike the latter, which handles all tasks simultaneously and relies on crowding distance for diversity, CycleQD alternates between tasks and naturally encourages diversity via the QD mechanism. This allows CycleQD to be more effective. Lastly, we suspect that the absence of a mechanism to rehabilitate temporarily underperforming models contributes to this discrepancy. Given the lengthy computation pipelines in LLMs, which include numerous layers, even a single malfunctioning layer can degrade overall performance. Unlike QD, the baselines do not have a way to recover these layers, potentially promising models are prematurely discarded.
@@ 4.1.5 Ablation studies
Sampling Methods: Finally, we demonstrate the importance of Elite sampling in CycleQD in trial 4, which has the identical settings as model 11 in Table 1, and gives the best score in the ablation studies. Elite sampling is effective because it prioritizes merging high-performing models, the result of which helps expand the Pareto frontier in each archive more effectively.
In summary, our ablation studies validate the importance of our design choices and shows that improvements aren’t merely additive - poor design choices can harm performance while our well-designed components work synergistically (e.g., comparing rows 1 and 2 in Table 2). The cumulative improvement from baseline QD to our final CycleQD is substantial at 4.8 percentage points (from 47.6% to 52.4%). This is a significant improvement in the context of LLMs performing complex agent tasks, particularly considering our model approaches gpt-3.5-turbo’s performance despite having significantly fewer parameters.
@@ 4.1.6 Skill generalization and language capabilities
In addition to agent skills, we evaluate our model on out-of-distribution coding tasks and language understanding across six categories, please see Appendix A.1.6 for detailed task descriptions. Table 3 shows the results across these tasks, where we normalize the scores against the Llama3-8B-Instruct base model. On average, CycleQD outperforms the three fine-tuned experts, providing better overall generalization across tasks. Our method generalizes well on the coding tasks while retraining performance on language tasks. In contrast, the MBPP expert, while delivering competitive scores on the coding tasks, shows significantly lower performance on specific tasks in the Reasoning category. This sharp decline in performance suggests the occurrence of catastrophic forgetting during fine-tuning, where the model loses its ability to generalize across tasks outside its specialization. These findings once again underscore the efficacy of CycleQD in achieving superior and more consistent performance across diverse tasks compared to traditional fine-tuning approaches.
@@ 4.2.1 Task description
Besides to its applications for LLMs, CycleQD serves as a versatile method for integrating expert models across various data modalities beyond text. For example, we include a vision question answering (VQA) task in addition to the CS tasks and find CycleQD to be able to outperform the experts (see Section A.1.5). In this experiment, we go further beyond and extend CycleQD to the fusion of multiple Segment Anything Models (SAM), which are state-of-the-art computer vision models designed for image segmentation tasks. Specifically, our objective is to merge pairs of SAM models, A and B, to create models whose capabilities encompass the skill sets of both A and B.
@@ 4.2.3 Results
5 Related works
LLM skill acquisition research has expanded beyond conversational tasks, focusing on enabling autonomous problem-solving through the LLM-as-Agent paradigm (Xi et al., 2023; Wang et al., 2024b). Techniques like ReAct (Yao et al., 2023) integrate reasoning with action, but skill acquisition through training remains limited, often relying on standard fine-tuning (Zeng et al., 2024; Qin et al., 2024), which struggles with balancing multiple skills. QD (Pugh et al., 2016) is an evolutionary paradigm aimed at discovering diverse, high-performing solutions. Algorithms like MAP-Elites (Mouret & Clune, 2015) and MOQD (Pierrot et al., 2022) have refined QD for tasks with diverse behaviors and conflicting objectives. While QD has been applied to neural networks and LLM prompts (Guo et al., 2024; Xue et al., 2024), evolving LLM weights remain unexplored. Model merging, which combines capabilities across pre-trained models at lower training costs, has advanced through techniques like sparsification (Yadav et al., 2023; Stoica et al., 2024) and optimization (Yang et al., 2024). Our work integrates QD with model merging, allowing LLM components to be preserved and optimized for diverse tasks. Please refer to Section B for more related works.
@@ Acknowledgements
The authors would like to thank Robert Lange for providing valuable discussions and feedback while drafting the text, and Kou Misaki and Qi Sun for providing infrastructure support. This paper is based on results obtained from a project, JPNP20017, subsidized by the New Energy and Industrial Technology Development Organization (NEDO).
@@ A.1.3 CycleQD hyper-parameters
Figure 5: Averaged performance on computer science tasks. Evaluated every 100 generations.
We set $\alpha_{\mathrm{low}}=0.5$ and $\alpha_{\mathrm{high}}=0.8$ in Elite sampling, $\mu=1.0$ and $\sigma=0.03$ in model merging-based crossover, and $w_{\text{max}}=0.3$ in our SVD-based mutations. These hyperparameters are determined in preliminary studies, where we arbitrarily determined a set of values, ran CycleQD for a small number of generations and picked the set with the best performance. We believe the performance of CycleQD can be further improved when incorporating sophisticated hyperparameter searching methods.
To investigate the influence of the number of generations on CycleQD performance, we analyzed the average scores of the computer science tasks every 100 generations. Figure 5 illustrates these results. The performance gradually increases over the generations with no obviously significant oscillations. While still showing a slight upward trend beyond this point, it stabilizes around 1000 generations.
@@
A.1.5 Adding a vision question answering task
In addition to the computer science tasks, we add a visual question answering (VQA) task to CycleQD to further demonstrate its generally applicability across tasks that differ drastically. The VQA task lies within the vision-language modeling (VLM) domain, which presents challenges when merging it with tasks from LLM domain.
Specifically, we use the TextVQA dataset (Singh et al., 2019) as part of the CycleQD experiments along with the coding, DB, and OS tasks. The experimental setup follows the same configurations described in Section 4.2.2, including parameters and the 1200 generations training budget. To create a VQA expert, we optimize the Llama3-Llava-Next-8B model using 4000 samples from the TextVQA data as training data. Additionally, we allocate 500 distinct samples, separated from the training data, for CycleQD training. These 500 samples are evenly divided into optimization and test datasets.
During the CycleQD optimization process, we extract the Llama3-8B-instruct component of the VLM and perform merge and mutate operations. The results, presented in Table 5, show that CycleQD achieves the best average performance compared to both the base and the expert models. Notably, CycleQD outperforms the several expert models on VQA, coding, and OS tasks, highlighting its effectiveness in combining knowledge across diverse domains.
Table 5: Evaluation on computer science tasks and VQA. We add VQA task to computer science tasks and evaluate CycleQD.
# Methods VQA MBPP DB OS Avg
0 Llama3-8B-Instruct (base model) 39.0 67.3 5.3 25.2 34.2
Experts models
1 VQA expert 51.4 4.0 0.0 1.5 14.2
2 Coding expert 32.9 70.4 21.2 20.7 36.3
3 DB expert 45.4 65.8 42.4 28.5 45.5
4 OS expert 46.1 66.3 0.0 30.4 35.7
5 CycleQD (Ours) 54.1 72.9 32.4 39.6 49.7